May 13-16

– Five stores provided ~63 hours of luminosity
– Tevatron personnel conducted hollow-beam studies
– EE support personnel helped operations turn on the Tevatron electron lens
– Booster suffered from low-level RF room temperature problems
– NuMI target problems held off NuMI beam
– Necessary cryo system repairs on wet engine flywheels required the run coordinator to terminate store 8737 on May 16

*The integrated luminosity for the period from 5/9/11 to 5/16/11 was 63.88 inverse picobarns. NuMI took 4.15E18 of beam during this same time period.

Read the Current Accelerator Update
Read the Early Bird Report
View the Tevatron Luminosity Charts